OpenStudy (anonymous):

In a parallelogram, one side equals 4y-13 and its opposite side equals 15-x. The other side equals 3x-5 and its opposite side equals y+1. Find x and y.

OpenStudy (paxpolaris):



OpenStudy (anonymous):

We can set up two equations: 4y - 13 = 15 - x 3x - 5 = y + 1 We can change the 2nd equation and solve for 'y': 3x - 5 = y + 1 Subtract 1 to both sides: 3x - 6 = y Plug in '3x - 6' for 'y' in the 1st equation: 4y - 13 = 15 - x 4(3x - 6) - 13 = 15 - x Distribute 4 into the parenthesis: 12x - 24 - 13 = 15 - x Combine like terms: 12x - 37 = 15 - x Add 'x' to both sides: 13x - 37 = 15 Add 37 to both sides: 13x = 52 Divide 13 to both sides: x = 4 Now we can plug this back into any of the two equations and solve for the 'y' value. 3x - 5 = y + 1 3(4) - 5 = y + 1 Multiply: 12 - 5 = y + 1 Combine: 7 = y + 1 Now subtract 1 to both sides, what's 7 - 1? @jackl

OpenStudy (anonymous):

Can you subtract 7 - 1? @jackl

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yes, so our full answer is: (4, 6)

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\sf\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!!!}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

I really appreciate it!! I understand now!

OpenStudy (anonymous):

No problem!